OpenStudy (anonymous):

There are 238 juniors at a school. The ratio of girls to boys is 3:4. How many girls are there? how many boys are there? Geometry Honors. Ratios and Propotions

OpenStudy (anonymous):

Ik how to do dis

OpenStudy (anonymous):

I just need somone to explain to me on how to get the answer, I have the answer but I don't know how to do the problem.

OpenStudy (studygurl14):

This is easy. Basically, all you have to do is split the amount of juniors into 7 parts. (3:4, 3+4 =7) 238 ÷ 7 = 34 Number of girls: 3/7, so 34 * 3 = 102 girls Boys: 4/7, 34*4 = 136 Think of it like fractions

OpenStudy (anonymous):

how'd you get 238? :P

OpenStudy (anonymous):

sounds bout right

OpenStudy (anonymous):

nvm lol

OpenStudy (anonymous):

Wow thanks I think I get it now.

OpenStudy (studygurl14):

Like I said. Think of it like fractions. Ratios are just another way of expressing fractions. 3:4, 3 girls for every 4 boys. If you have 7 people, 3 are girls, 4 are boys. 3/7 girls, 4/7 boys If you have 238 people, and 3/7 are girls, and 4/7 are boys....what do you do to find how many boys and how many girls?